OpenStudy (anonymous):

The sum of the squares of two consecutive positive even numbers is 52. What is the smaller number? @tcarroll010

OpenStudy (anonymous):

Let be a and b the two consecutive positive even numbers. So : \[a=2n~~~ b=a+2=2n+2\] And then : \[a^2+b^2=52\\\] Then : \[4n^2+(2n+2)^2=52\] So : \[4n^2+4n^2+8n+4=52\] So : \[8n^2+8n-48=0\] We divide by 8 : \[n^2+n-6=0\] SO : \[\Delta=1-4\times(-6)+25\] So : \[n=\frac{-1+\sqrt{25}}{2}=2\] So the smallest number is : \[a=2n=2\times2=4\]